OpenStudy (anonymous):

Algebra Help, please, I'm desperate! Given the function f(x) = 3x, find the value of f^−1(81).

OpenStudy (johnweldon1993):

That means to find the "inverse" of this function

OpenStudy (johnweldon1993):

All you need to do is switch the positions of the 'x' and 'y' in your original equation...and then use algebra to solve for 'y' again

OpenStudy (johnweldon1993):

\[\large y = 3x\] Switch the 'x' and 'y' \[\large x = 3y\] How would you solve for 'y' again?

OpenStudy (anonymous):

Step 1: Replace f(x) with y. Step 2: Divide both sides by 3. Step 3: Switch y and x. Step 4: Replace y with \(f^{-1}(x)\) Step 5: Plug in 81 and solve.

OpenStudy (anonymous):

Oh, sorry, it's f^-1(x)=3^x, and I got that far but I don't know how to solve if x=3^y

OpenStudy (anonymous):

Take the log of both sides?

OpenStudy (johnweldon1993):

Ahh okay... \[\large y = 3^x\] \[\large x = 3^y\] Take the log of both sides since \(\large lna^x = x \times ln(a)\) So we have \[\large ln(x) = y \times ln(3)\]

OpenStudy (anonymous):

so y=1? Meaning f^-1=1?

OpenStudy (anonymous):

You still need to isolate y on the right-side. So \[\frac{lnx}{ln3} = y\]

OpenStudy (anonymous):

And then you need to replace y with \(f^{-1}(x)\), and then plug in 81.

OpenStudy (johnweldon1993):

Indeed^ You forgot that step @1234k123a12l1i so \[\large y = \frac{ln(x)}{ln(3)}\] Now plug in x = 81

OpenStudy (anonymous):

\[\log_{81} \div \log_{3} = f^-1\]

OpenStudy (anonymous):

Can I divide the 81 and the 3 both by 3?

OpenStudy (anonymous):

Keep in mind (not important to the problem, but for understanding), \[\frac{lnx}{ln3} = \log_3(x)\] and @johnweldon1993 showed you how to enter that into your calculator correctly :).

OpenStudy (anonymous):

Or I am imagining things....

OpenStudy (johnweldon1993):

You are correct @meshlogic :)

OpenStudy (anonymous):

Ohh okay I get it now, thank you so much you guys are awesome!

OpenStudy (anonymous):

Plug this into your calc: \[\frac{log81}{log3} = log81 \div log3\]

OpenStudy (anonymous):

Yeah I just did that, and got 4

OpenStudy (johnweldon1993):

So basically \[\large log_3(x)\] Plug in 81 for 'x' \[\large log_3(81) = x\] It is asking what power of 3 do you need to make 81? \[\large 3^x = 81\]

OpenStudy (johnweldon1993):

Which yes IS 4 :)

OpenStudy (anonymous):

Great :D That's right!

OpenStudy (anonymous):

What a team. Have a nice day. :)

OpenStudy (johnweldon1993):

You as well @meshlogic :)

OpenStudy (anonymous):

Thank you, you too :)